OpenStudy (anonymous):

What is the best answer to this problem? 3x(2x-1) + (2x-1)=

OpenStudy (***[isuru]***):

Hi, u can take out (2x -1) and factorize it again... \[3x( 2x - 1) + (2x - 1) = (2x - 1)(3x +1)\] and if the expression is equal to zero...then u can solve it for "x" where u will get the answers like this.. \[(2x - 1)(3x +1) =0\]\[(2x -1) = 0 \ or \ (3x+ 1) =0\]\[x = \frac{ 1 }{2 } \ or \ x = -\frac{ 1 }{ 3 }\] Hope this will help ya !!!